# Ground-based Image Deconvolution with Swin Transformer UNet

### v1

@@ Abstract
Aims. As ground-based all-sky astronomical surveys will gather millions of images in the coming years, a critical requirement emerges for the development of fast deconvolution algorithms capable of efficiently improving the spatial resolution of these images. By successfully recovering clean and high-resolution images from these surveys, our objective is to help deepen our understanding of galaxy formation and evolution through accurate photometric measurements.
@@
Results. By conducting a performance comparison between our deep learning-based method and Firedec, a classical deconvolution algorithm, we analyze a subset of the EDisCS cluster samples. We demonstrate the advantage of our method in terms of resolution recovery, generalization to different noise properties, and computational efficiency. Not only does the analysis of this cluster sample assess the efficiency of our method, but it also enables us to quantify the number of clumps within these galaxies in relation to their disc colour. This robust technique holds promise for identifying structures in the distant universe from ground-based images.
@@
High spatial resolution and high signal-to-noise observations are prerequisites to most observational astrophysical problems. However, expecting the two conditions to happen simultaneously is challenging: space telescopes have limited collecting power, while large telescopes are ground-based and are therefore affected by atmospheric turbulence. This is clearly illustrated by the two main missions/observatories of the next decade: the ESA-NASA Euclid space telescope (Euclid Collaboration et al., 2022; Laureijs et al., 2011) and the Vera C. Rubin Observatory (Ivezić et al., 2019). Exploiting the best of both worlds is possible, provided reliable post-processing techniques are developed to remove blurring by the atmosphere and instrument Point Spread Function (PSF). To further complicate matters, sensor variations introduce noise in the images. Hence, image deconvolution in astrophysics is an ill-posed and ill-conditioned inverse problem that requires regularization to achieve a unique solution. This was realized very early in the field, and solutions were proposed, such as minimizing the Tikhonov function (e.g. Tikhonov & Arsenin, 1977) or maximizing the entropy of the solution (e.g. Skilling & Bryan, 1984). Other algorithms, based on Bayesian statistics, include the Lucy-Richardson algorithm, used on the early data from the Hubble Space Telescope (HST) (Richardson, 1972; Lucy, 1974). Magain et al. (MCS; 1998) proposed a two-channel method that separates the point sources from the spatially extended ones and that deconvolves using a narrow PSF, hence achieving finite but improved resolution well compatible with the sampling (pixel) chosen to represent the solution. Improvements of the ”MCS” method implement wavelet regularization of the extended channel of the solution Cantale et al. (2016a). This method was further refined by Michalewicz et al. (STARRED; 2023) who use an isotropic wavelet basis, called Starlets (Starck et al., 2015), to regularize the solution.
@@
However, deep learning offers a completely different approach, by learning the properties of the desired solution. Once trained, deep learning-based methods are also orders of magnitude faster than classical methods. Of note, UNets (Ronneberger et al., 2015) have become popular due to their highly non-linear processing and multi-scale approach. Building on Unet, Sureau et al. (2020) developed the Tikhonet method for deconvolving galaxy images in the optical domain. They demonstrated that Tikhonet outperformed sparse regularization-based methods in terms of Mean Squared Error (MSE) and a shape criterion, where a measure of the galaxy ellipticity was used to encode its shape. Nammour et al. (2022) added a shape constraint to the Tikhonet loss function and achieved better performance. In our recent work (Akhaury et al., 2022), we proposed a new deconvolution approach that employs the Learnlet decomposition (Ramzi et al., 2023), using the same two-step approach as in Sureau et al. (2020), but by substituting the Unet denoiser by Learnlet.
@@
With the recent advent of Vision Transformers (Dosovitskiy et al., 2021), significant progress has been made in the field of image restoration (Liang et al., 2021; Zamir et al., 2022; Wang et al., 2022). This is the motivation for the present work to investigate the performance of SUNet (Fan et al., 2022), a variant of Unet with Swin Transformer blocks (Liu et al., 2021) replacing the convolutional layers. To our knowledge, SUNet has not yet been used as a denoiser in a deconvolution framework. We show that the neural network outputs introduce bias, thereby limiting the scientific analysis. The bias appears as a small flux loss in the outputs. To counter this, we propose a third debiasing procedure based on the active coefficients in the sparsity wavelet framework, called Multi-resolution support (Starck et al., 1995), as explained in Section 2.3. Our experiments involved real HST images, with the network trained on images extracted from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011).
@@
Finally, to assess the generalizability, we test our new code on a completely different sample of ground-based images obtained with the FORS2 camera at the Very Large Telescope (VLT). Spatially resolved images of galaxies can help address a multitude of topics involving their morphologies. In the present case, we tackle the question of how the galaxy cluster environment could impact the properties of the discs of spiral galaxies. From the same dataset, Cantale et al. (2016b) had found that of cluster spiral galaxies at redshift - had disc colours redder by more than of the mean colours of their field counterparts. This result was obtained thanks to the deconvolution code Firedec (Cantale et al., 2016a). The VLT and -band images, with initial spatial resolutions between and had been deconvolved with a target final resolution of on pixels. Even though the gain in resolution was indeed substantial, conditions were not sufficient to go beyond the global photometric properties of the discs. In this study, we go one step further and investigate their internal structure, specifically by identifying their star-forming clumps.
@@
Compact star-forming clumps are identified in distant galaxies, particularly when they benefit from HST deep images (e.g., Wuyts et al., 2012; Guo et al., 2015; Sattari et al., 2023). They are understood as playing a crucial role in galaxy assembly and star-formation activity. Recent research by Sok et al. (2022) explored clump fractions in star-forming galaxies (SFGs) from multi-band analysis in the COSMOS field. The match between HST and ground-based resolution was performed with Firedec. Their findings indicated a decline in the fraction of clumpy galaxies with increasing stellar masses and redshifts. Moreover, they observed that clumps contributed a higher fractional mass towards galaxies at higher redshifts. In our study, by employing a more powerful deconvolution algorithm capable of accurately recovering small-scale structures at high spatial resolution from ground-based multi-band observations (as demonstrated in Section 4.2), our goal is to quantify the number of clumps in EDisCS cluster galaxies and examine their relationship with disc colour.
@@
In Section 2, we present the deconvolution problem and introduce our proposed deep learning method to address it. The process of generating our dataset and conducting experiments is outlined in Section 3. In Section 4, we present the results of our deconvolution algorithm. Finally, we draw our conclusions in Section 5. To support reproducible research, the codes and trained models utilized in this article are publicly accessible in Section 6. Additional studies and supplementary information can be found in Appendices A, B, and C.
2 Deep-Learning based Deconvolution
@@
The deconvolution problem can be summarized in a very simple (but hard to solve) equation. Let be the observed image and be the PSF. The observed image can be modelled as
@@
where denotes the target image, denotes the convolution operation, and denotes additive Gaussian noise. The goal is then to recover the ground truth image , given the PSF convolution and the unknown noise. Such ill-posed inverse problems require regularization of the solution, to select the most appropriate one among the many others that are compatible with the data. Sparse wavelet regularization using the or norm has remained the most accepted one in the past, but the recent advent of machine learning methods has changed this paradigm.
@@ 2.1 Tikhonov Deconvolution
Tikhonov Deconvolution is a two-step deep learning-based deconvolution technique. In the first step, the input images undergo deconvolution using a Tikhonov filter with quadratic regularization. If denotes the circulant matrix associated with the convolution operator , the Tikhonov solution of equation 1 is expressed as:
Here, represents the linear Tikhonov filter, configured as a Laplacian high-pass filter to penalize high frequencies. The regularization weight, denoted as , is determined through a grid search. The Tikhonov step produces deconvolved images containing correlated additive noise, which is subsequently eliminated in the following step by an appropriate denoiser. The denoisers are trained to learn the mapping from the Tikhonov output to the ground truth image by minimizing a suitable loss function, such as or .
The denoising performance is significantly influenced by the choice of the model architecture. To effectively capture distant correlations, it is crucial to incorporate multi-scale processing in the model design. This consideration leads to the adoption of a layout similar to that of a UNet (Ronneberger et al., 2015). Additionally, Mohan et al. (2020) demonstrated that biases in convolutional layers can result in low generalization capability. Consequently, for our experiments, we opt for bias-free networks.
@@ 2.2 SUNet Denoising
In recent times, the UNet architecture has become popular in various image-processing applications due to its incorporation of hierarchical feature maps, facilitating the acquisition of multi-scale contextual features. It is widely employed in diverse computer vision tasks, including segmentation and restoration (Yu et al., 2019; Gurrola-Ramos et al., 2021). Evolved versions such as Dense-UNet (Guan et al., 2020), Res-UNet (nan Xiao et al., 2018), Non-local UNet (Yan et al., 2020), and Attention UNet (Jin et al., 2020) have also emerged. Thanks to its flexible structure, the UNet can adapt to different building blocks, enhancing its overall performance. Moreover, the evolution of image processing methodologies has seen the introduction of Transformer models (Vaswani et al., 2017), which were initially successful in natural language processing (NLP) but have also demonstrated impressive performance in image classification (Dosovitskiy et al., 2021; Yuan et al., 2021). However, when directly applied to vision tasks, transformers face challenges such as the large-scale difference between images and sequences, making them less effective in modelling long sequences due to the need for a square number of parameters for 1-dimensional sequences. Additionally, transformers are not well-suited for dense prediction tasks such as instance segmentation at the pixel level. Swin Transformer addresses these issues by introducing a shifted-window mechanism to reduce the number of parameters, establishing itself as a state-of-the-art solution for high-level vision tasks (Liu et al., 2021). Taking inspiration from these advancements, Fan et al. (2022) incorporated Swin Transformers as building blocks within the UNet architecture and showed that they could achieve competitive results compared to existing benchmarks for image denoising. The network architecture is visually depicted in Figure 1, and the PyTorch code is publicly available on GitHub (details in Section 6).
While SUNet was originally developed for White Gaussian Noise (WGN) removal, the Tikhonov deconvolution step (equation 2) introduces alterations to the noise characteristics in the images. Consequently, it becomes crucial to assess the extent to which SUNet can effectively handle the presence of Colored Gaussian Noise (CGN). Thus, our tests examine SUNet’s generalizability in the presence of CGN.
@@ 2.3 Debiasing with Multi-resolution Support
Our study reveals that the deep learning-based solution introduces a bias, evident in the form of positive structures in the residuals, as illustrated in Figure 2(a). This bias can have implications for the accuracy of scientific analyses, potentially influencing the flux estimation within the recovered features in the reconstructed images. To recover the lost flux and enhance image sharpness, multi-resolution support has proven effective (Starck et al., 1995). Denoting the Starlet transform as and the SUNet output solution as , an iterative process allows the recovery of flux from the residual . In each iteration, denoted as , a debiasing correction term is applied by multiplying the multi-resolution support matrices of the SUNet output solution at each scale with the Starlet decomposition of the gradient of the residual, incrementally modifying the deconvolved image:
@@
The acronym MRS stands for multi-resolution support, indicating a boolean measure of whether an image contains information at a specific pixel and scale. If represents the wavelet coefficient at a given scale and is the threshold value, the multi-resolution support operation can be defined as:
The flux-recovery process is illustrated in Figure 2. The iterative process is stopped once convergence is achieved in the standard deviation of the residual as a function of the number of iterations, as seen in sub-figure 2(d). A more detailed study on the impact of debiasing with multi-resolution support on neural networks is presented in Appendix B.
@@ 3.1 Training Dataset Generation
We extracted Hubble Space Telescope (HST) cutouts measuring pixels from CANDELS (Grogin et al., 2011; Koekemoer et al., 2011) in the filter (-band). These cutouts were then convolved with a Gaussian PSF having a Full Width at Half Maximum (FWHM) of 15 pixels. Following the convolution, we injected White Gaussian Noise (WGN) with a standard deviation denoted as . This choice ensured that the faintest object in our dataset had a peak Signal-to-Noise Ratio (SNR) close to and was barely visible. With this particular value of , our dataset exhibited a range of SNR values depending on the magnitude of each galaxy. To standardize the images, each one was normalized within the range. This normalization involved subtracting its own mean and scaling its peak value by the maximum value in the image. Finally, the batch of images was randomly divided into Training-Validation-Test subsets in the ratio .
@@ 3.2 Training the SUNet
The SUNet architecture was trained using a Titan RTX Turing GPU with 24 GB RAM for each job. The training aimed at learning the mapping from the Tikhonov output (containing CGN) to the corresponding HST image . The training utilized the Adam optimizer (Kingma & Ba, 2014) with an initial learning rate of , gradually halving every 25 epochs until reaching a minimum of . As in the original SUNet paper (Fan et al., 2022), the -loss was used for training. A more detailed discussion on the significance of the training loss function is given in Appendix C. The input images were processed in mini-batches of size . The dataset was augmented with random rotations in multiples of 90°, translations and flips along horizontal & vertical axes. Starting with images in our initial training dataset, we increase its diversity by a factor of by performing augmentation.

### v2

@@ Abstract
Aims. As ground-based all-sky astronomical surveys will gather millions of images in the coming years, a critical requirement emerges for the development of fast deconvolution algorithms capable of efficiently improving the spatial resolution of these images. By successfully recovering clean and high-resolution images from these surveys, the objective is to deepen the understanding of galaxy formation and evolution through accurate photometric measurements.
@@
Results. We conducted a performance comparison between our deep learning-based method and Firedec, a classical deconvolution algorithm, based on an analysis of a subset of the EDisCS cluster samples. We demonstrate the advantage of our method in terms of resolution recovery, generalisation to different noise properties, and computational efficiency. The analysis of this cluster sample not only allowed us to assess the efficiency of our method, but it also enabled us to quantify the number of clumps within these galaxies in relation to their disc colour. This robust technique that we propose holds promise for identifying structures in the distant universe through ground-based images.
@@
High spatial resolution and high signal-to-noise observations are prerequisites to most observational astrophysical problems. However, expecting the two conditions to happen simultaneously is challenging, as space telescopes have a limited collecting power, while large telescopes are ground-based and are therefore affected by atmospheric turbulence. This is clearly illustrated by the two main missions of the next decade: the ESA-NASA Euclid space telescope (Euclid Collaboration et al., 2022; Laureijs et al., 2011) and the Vera C. Rubin Observatory (Ivezić et al., 2019). Exploiting the best of both worlds is possible, provided reliable post-processing techniques are developed to remove blurring by the atmosphere and instrument point spread function (PSF). To further complicate matters, sensor variations introduce noise into images. Hence, image deconvolution in astrophysics is an ill-posed and ill-conditioned inverse problem that requires regularisation to achieve a unique solution. This was realised very early in the field, and solutions were proposed, such as minimising the Tikhonov function (e.g. Tikhonov & Arsenin, 1977) or maximizing the entropy of the solution (e.g. Skilling & Bryan, 1984). Other algorithms, based on Bayesian statistics, include the Lucy-Richardson algorithm, used on the early data from the Hubble Space Telescope (HST; Richardson (1972); Lucy (1974)). Magain et al. (MCS; 1998) proposed a two-channel method that separates the point sources from the spatially extended ones and deconvolves using a narrow PSF, hence achieving finite but improved resolution compatible with the sampling (pixel) chosen to represent the solution. Improvements of the ‘MCS’ method implement wavelet regularisation of the extended channel of the solution Cantale et al. (2016a), and this improved method was further refined by Michalewicz et al. (STARRED; 2023), who used an isotropic wavelet basis called Starlets (Starck et al., 2015) to regularise the solution.
@@
However, deep learning offers a completely different approach, by learning the properties of the desired solution. Once trained, deep learning-based methods are also orders of magnitude faster than classical methods. Of note, UNets (Ronneberger et al., 2015) have become popular due to their highly non-linear processing and multi-scale approach. Building on Unet, Sureau et al. (2020) developed the Tikhonet method for deconvolving galaxy images in the optical domain. They demonstrated that Tikhonet outperformed sparse regularisation-based methods in terms of mean squared error and a shape criterion, where a measure of the galaxy ellipticity was used to encode its shape. Nammour et al. (2022) added a shape constraint to the Tikhonet loss function and achieved better performance. In our recent work (Akhaury et al., 2022), we have proposed a new deconvolution approach that employs the Learnlet decomposition (Ramzi et al., 2023). It uses the same two-step approach as in Sureau et al. (2020) but substitutes the UNet denoiser by Learnlet.
@@
With the recent advent of Vision Transformers (Dosovitskiy et al., 2021), significant progress has been made in the field of image restoration (Liang et al., 2021; Zamir et al., 2022; Wang et al., 2022). This is the motivation for the present work to investigate the performance of SUNet (Fan et al., 2022), a variant of Unet with Swin Transformer blocks (Liu et al., 2021) replacing the convolutional layers. To our knowledge, SUNet has not yet been used as a denoiser in a deconvolution framework. We show that the neural network outputs introduce bias, thereby limiting the scientific analysis. The bias appears as a small flux loss in the outputs. To counter this, we propose a third debiasing procedure based on the active coefficients in the sparsity wavelet framework called multi-resolution support (Starck et al., 1995), explained in Section 2.3. Our experiments involved real HST images, and the network was trained on images extracted from the CANDELS survey (Grogin et al., 2011; Koekemoer et al., 2011).
@@
Finally, to assess its generalisability, we tested our new code on a completely different sample of ground-based images obtained with the FORS2 camera at the Very Large Telescope (VLT). Spatially resolved images of galaxies can help address a multitude of topics involving their morphologies. In the present case, we tackle the question of how the galaxy cluster environment can impact the properties of the discs of spiral galaxies. From the same dataset, Cantale et al. (2016b) found that of cluster spiral galaxies at redshift - have disc colours that are redder by more than of the mean colours of their field counterparts. This result was obtained thanks to the deconvolution code Firedec (Cantale et al., 2016a). The VLT - and -band images, with initial spatial resolutions between and , were deconvolved with a target final resolution of on pixels. Even though the gain in resolution was indeed substantial, conditions were not sufficient to go beyond the global photometric properties of the discs. In this study, we go one step further and investigate their internal structure, specifically by identifying their star-forming clumps.
@@
Compact star-forming clumps have been identified in distant galaxies, particularly with the aid of HST deep images (e.g., Wuyts et al., 2012; Guo et al., 2015; Sattari et al., 2023). They are understood to play a crucial role in galaxy assembly and star-formation activity. Recent research by Sok et al. (2022) explored clump fractions in star-forming galaxies from multi-band analysis in the COSMOS field. The match between HST and ground-based resolution was performed with Firedec. Their findings indicated a decline in the fraction of clumpy galaxies with increasing stellar masses and redshifts. Moreover, they observed that clumps contributed a higher fractional mass towards galaxies at higher redshifts. In our study, by employing a more powerful deconvolution algorithm capable of accurately recovering small-scale structures at high spatial resolution from ground-based multi-band observations (as demonstrated in Section 4.2), our goal is to quantify the number of clumps in EDisCS cluster galaxies and examine their relationship with disc colour.
@@
In Section 2, we present the deconvolution problem and introduce our proposed deep learning method to address it. The process of generating our dataset and conducting experiments is outlined in Section 3. In Section 4, we present the results of our deconvolution algorithm. Finally, we draw our conclusions in Section 5. To support reproducible research, the codes and trained models utilised in this article are publicly accessible in Section 6. Additional studies and supplementary information can be found in Appendices A, B, and C.
2 Deep learning-based deconvolution
@@
The deconvolution problem can be summarised with a very simple (but hard to solve) equation. Let be the observed image and be the PSF. The observed image can be modelled as
@@
where denotes the target image, denotes the convolution operation, and denotes additive Gaussian noise. The goal is to recover the ground truth image , given the PSF convolution and the unknown noise. Such ill-posed inverse problems require regularisation of the solution in order to select the one that is most appropriate compared to the many others that are compatible with the data. Sparse wavelet regularisation using the or norm remained the most accepted regularisation in the past, but the recent advent of machine learning methods has changed the paradigm.
@@ 2.1 Tikhonov deconvolution
Tikhonov Deconvolution is a two-step deep learning-based deconvolution technique. In the first step, the input images undergo deconvolution using a Tikhonov filter with quadratic regularisation. If denotes the circulant matrix associated with the convolution operator , the Tikhonov solution of equation 1 is expressed as
Here, represents the linear Tikhonov filter, configured as a Laplacian high-pass filter to penalise high frequencies. The regularisation weight, denoted as , is determined through a grid search. The Tikhonov step produces deconvolved images containing correlated additive noise, which is subsequently eliminated in the following step by an appropriate denoiser. The denoisers are trained to learn the mapping from the Tikhonov output to the ground truth image by minimising a suitable loss function, such as or .
The denoising performance is significantly influenced by the choice of the model architecture. To effectively capture distant correlations, it is crucial to incorporate multi-scale processing in the model design. This consideration leads to the adoption of a layout similar to that of a UNet (Ronneberger et al., 2015). Additionally, Mohan et al. (2020) demonstrated that biases in convolutional layers can result in a low generalisation capability. Consequently, for our experiments, we opted for bias-free networks.
@@ 2.2 SUNet denoising
In recent times, the UNet architecture has become popular in various image-processing applications due to its incorporation of hierarchical feature maps, which facilitate the acquisition of multi-scale contextual features. It is widely employed in diverse computer vision tasks, including segmentation and restoration (Yu et al., 2019; Gurrola-Ramos et al., 2021). Evolved versions such as Dense-UNet (Guan et al., 2020), Res-UNet (nan Xiao et al., 2018), Non-local UNet (Yan et al., 2020), and Attention UNet (Jin et al., 2020) have also emerged. Thanks to its flexible structure, UNet can adapt to different building blocks, enhancing its overall performance. Moreover, the evolution of image-processing methodologies has seen the introduction of transformer models (Vaswani et al., 2017), which were initially successful in natural language processing but have also demonstrated impressive performance in image classification (Dosovitskiy et al., 2021; Yuan et al., 2021). However, when directly applied to vision tasks, transformers face challenges such as the large-scale difference between images and sequences, making them less effective in modelling long sequences due to a need for a square number of parameters for one-dimensional sequences. Additionally, transformers are not well suited for dense prediction tasks such as instance segmentation at the pixel level. Swin Transformer addresses these issues by introducing a shifted-window mechanism to reduce the number of parameters, establishing itself as a state-of-the-art solution for high-level vision tasks (Liu et al., 2021). Taking inspiration from these advancements, Fan et al. (2022) incorporated Swin Transformers as building blocks within the UNet architecture and showed that they could achieve competitive results compared to existing benchmarks for image denoising. The network architecture is visually depicted in Figure 1, and the PyTorch code is publicly available on GitHub (details in Section 6).
While SUNet was originally developed for white Gaussian noise removal, the Tikhonov deconvolution step (equation 2) introduces alterations to the image’s noise characteristics. Consequently, it becomes crucial to assess the extent to which SUNet can effectively handle the presence of coloured Gaussian noise. Thus, our tests examine SUNet’s generalisability in the presence of CGN.
@@ 2.3 Debiasing with multi-resolution support
Our study reveals that the deep learning-based solution introduces a bias, evident in the form of positive structures in the residuals, as illustrated in Figure 2(a). This bias can have implications on the accuracy of the scientific analyses, potentially influencing the flux estimation within the recovered features in the reconstructed images. In recovering the lost flux and enhancing image sharpness, multi-resolution support has been proven effective (Starck et al., 1995). Denoting the Starlet transform as and the SUNet output solution as , an iterative process allowed for the recovery of flux from the residual . In each iteration, denoted as , a debiasing correction term was applied by multiplying the multi-resolution support matrices of the SUNet output solution at each scale with the Starlet decomposition of the gradient of the residual, incrementally modifying the deconvolved image:
@@
, and
The acronym ‘MRS’ stands for multi-resolution support and indicates a Boolean measure of whether an image contains information at a specific pixel and scale. If represents the wavelet coefficient at a given scale and is the threshold value, the multi-resolution support operation can be defined as:
The flux-recovery process is illustrated in Figure 2. The iterative process was stopped once convergence was achieved in the standard deviation of the residual as a function of the number of iterations, as seen in sub-figure 2(d). A more detailed study on the impact of debiasing with multi-resolution support on neural networks is presented in Appendix B.
@@ 3.1 Training dataset generation
We extracted HST cutouts measuring pixels from CANDELS (Grogin et al., 2011; Koekemoer et al., 2011) in the filter (-band). These cutouts were then convolved with a Gaussian PSF having a full width at half maximum (FWHM) of 15 pixels. Following the convolution, we injected white Gaussian noise with a standard deviation denoted as . This choice ensured that the faintest object in our dataset had a peak signal-to-noise (S/N) close to one and was barely visible. With this particular value of , our dataset exhibited a range of S/N values depending on the magnitude of each galaxy. To standardise the images, each image was normalised within the range. This normalization involved subtracting the image’s mean and scaling the peak value by the image’s maximum value. Finally, the batch of images was randomly divided into training-validation-test subsets in the ratio .
@@ 3.2 Training the SUNet
The SUNet architecture was trained using a Titan RTX Turing GPU with 24 GB RAM for each job. The training aimed at learning the mapping from the Tikhonov output (containing CGN) to the corresponding HST image . The training utilised the Adam optimiser (Kingma & Ba, 2014) with an initial learning rate of , gradually halving every 25 epochs until reaching a minimum of . As in the original SUNet paper (Fan et al., 2022), -loss was used for training. A more detailed discussion on the significance of the training loss function is given in Appendix C. The input images were processed in mini-batches of size . The dataset was augmented with random rotations in multiples of 90°; translations and flips were along horizontal and vertical axes. Starting with images in our initial training dataset, we increased its diversity by a factor of ten by performing augmentation.